OpenStudy (anonymous):

how to convert -450° from degrees to radians.

OpenStudy (jdoe0001):

how many radians in \(180^o\)?

OpenStudy (jdoe0001):

anyhow... look at your Unit Circle there's one \(\pi\) radians\ in \(180^o\)

OpenStudy (anonymous):

I actually do not know, im looking at the unit circle right now

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} radians&degrees \\\hline\\ \pi &180\\ x&-450 \end{array}\implies \cfrac{\pi }{x}=\cfrac{180}{-450}\)

OpenStudy (jdoe0001):

so, just solve for "x"

OpenStudy (anonymous):

okay so it would be -5pi/2?

OpenStudy (jdoe0001):

yeap